OpenStudy (anonymous):

What is the y-intercept for this linear equation? 1/2x-2/3y=-6. *Straight up answer would be nice.*

OpenStudy (okra77):

multiply everything 6 to get rid of the fractions then solve for y the y intercept is 0,9

OpenStudy (anonymous):

Substitute 0 for x & solve for y 1/2(0) - 2/3y = -6 - 2/3y = -6 Multiply by -3/2 on each side of eq y = -9 Does that help?

OpenStudy (anonymous):

Y=9

OpenStudy (anonymous):

Lost my negative. The y intercept is (0,9)